Given the equation y= 1/2(x -5)^2 - 3
A write the equation in standard form
b Write the equation in intercept form
c find the x intercept(s) and the y intercept or explanin why the intercepts do not exist
d. graph using the given form.
I really just need help with part b

Answer by stanbon(75887)

y= 1/2(x -5)^2 - 3
A write the equation in standard form
b Write the equation in intercept form
---
Form:: y = a(x-b)(x-c)
----------------------
y = (1/2)(x^2-10x+25) - 3
I graphed it and found x-intercepts at x = 2.55 and 7.45
----
y = a(x-2.55)(x-7.45)
----
Solve for "a" using the y-intercept
y-intercept:
y = (1/2)(0-5)^2-3
y = 12.5-3 = 9.5
------
y-intercept = (0,9.5)
---
Solve for "a":
9.5 = a(-2.55)(-7.45)
a = 0.5
---
Intercept Form::
y = 0.5(x-2.55)(x-7.45)
----------------
Cheers,
Stan H.
---------------
c find the x intercept(s) and the y intercept or explainin why the intercepts do not exist
d. graph using the given form.
I really just need help with part b